All of MER-9 (AC23,24,25,26-9) has now moved to NAE-11 trunk A. All old mappings have been deleted and the user views are corrected 🙂
Due to burnt popcorn I have overridden the mixed air damper on 5305’s AHU 1 serving the board room and neighbooring offices. Please release later today or tonight.
2;50 PM update: I have also overridden 5305 B unit. Units A & C were already running 100% o.a. at this time. The smell is now affecting corridors as far as AP Lab. Awful!
Troubleshoot & found failed damper actuator. Replaced VMA (VMA-0384) & downloaded program. Rechecked all okay.
Worked with Fredasys and Kyle from JCI to improve METASYS response related to OSC chilled water and AHU 53-32’s d.a.t. set points. It was noted by OSC staff that the procedure room’s humidity levels seemed high; although within our allowable range (20%-60%), it was determined that a better reset schedule that includes r.a. enthalpy would enhance the system’s performance. Dr. Mazzeo should be happier…
Met with Kyle from JCI (new construction warranty) today to review nuisance alarms/system operation related to the OSC. He will eliminate the high disharge air humidity alarms (Thank you Cori for letting us know) as well tightening up the performance of O.R. heating valves etc. response times (Thank you Fredasys).